### v2

# Unified Unbiased Variance Estimation for Maximum Mean Discrepancy: Robust Finite-Sample Performance with Imbalanced Data and Exact Acceleration under Null and Alternative Hypotheses

Abstract
The maximum mean discrepancy (MMD) is a kernel-based nonparametric statistic for two-sample testing, whose inferential accuracy depends critically on variance characterization. Existing work provides various finite-sample estimators of the MMD variance, often differing under the null and alternative hypotheses and across balanced or imbalanced sampling schemes. In this paper, we study the variance of the MMD statistic through its U-statistic representation and Hoeffding decomposition, and establish a unified finite-sample characterization covering different hypotheses and sample configurations. Building on this analysis, we propose an exact acceleration method for the univariate case under the Laplacian kernel, which reduces the overall computational complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$.
1 Introduction
The fundamental problem of nonparametric two-sample testing is to detect differences between two probability distributions given access only to samples drawn from them. Maximum Mean Discrepancy (MMD) is a kernel-based nonparametric statistic proposed for this task [6, 11]. When equipped with a characteristic kernel, MMD defines a metric on the space of probability distributions, and thus equals zero if and only if the two distributions are identical [22, 27]. It transforms the distance measure between any two distributions into a linear operation in the Hilbert space defined by a kernel function, which greatly facilitates computations. For this reason, this method has attracted increasing attention in recent years. In particular, MMD has been successfully applied to tasks such as domain adaptation[23] and generative modeling[29], and has shown promising potential in other areas, such as Bayesian inference and model calibration tasks [31, 32].
The inferential reliability of MMD, particularly in constructing hypothesis tests and confidence intervals, depends crucially on an accurate characterization of its variance. However, the asymptotic behavior of the unbiased MMD statistic is non-trivial, leading to distinct variance structures under different regimes. Under the null hypothesis, the statistic is a degenerate U-statistic, with a limiting distribution given by an infinite weighted sum of chi-square random variables. Under the alternative hypothesis, it converges to a Gaussian distribution [11, 3]. This dichotomy makes variance estimation challenging, especially in practice where sample sizes from the two distributions often differ. Existing work has addressed various special cases: Gretton et al. [11] provided the unbiased variance estimator under the null hypothesis with equal sample sizes; Sutherland and Deka [28] characterized the unbiased MMD estimator as a U-statistic under balanced sampling; and Wei et al. [30] proposed a method under the alternative hypothesis with unequal sample sizes by treating the MMD estimator as a unified generalized U-statistic [18, 25], although their variance estimator is based on asymptotic approximations and remains biased in finite-sample scenarios. In this context, a unified formulation capable of providing unbiased variance estimation under both hypotheses while accommodating unequal sample sizes remains of significant theoretical and practical interest.
To address these limitations, our approach employs the Hoeffding decomposition to explicitly construct orthogonal projection terms [26, 14], providing a streamlined representation of the MMD variance estimator. This formulation implicitly incorporates the covariance components present in [30], while simplifying the overall structure. The orthogonal structure further enables a unified and smooth modeling of the variance transition across the null and alternative hypotheses. Beyond its theoretical elegance, the unbiased estimator’s algebraic clarity facilitates a concise, tensorized implementation in modern deep learning frameworks [1, 24], offering a practical and interpretable objective for kernel optimization aimed at maximizing test power [29, 13, 21, 7].
Beyond the theoretical perspective, accurately estimating the MMD variance poses significant computational challenges, as naive summation incurs quadratic complexity, straining both memory and computational resources and limiting the practical deployment of recent theoretical advances. Existing acceleration strategies exhibit clear trade-offs: global approximation methods, such as subsampling or random feature projections, scale well to high-dimensional data but introduce inherent approximation errors [33, 34]. Analytical acceleration offers an alternative, restructuring independent summations into ordered linear recursions [2, 17], a strategy previously applied in classical statistics for measures like Kendall’s tau and distance covariance to reduce complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$ [19, 16]. In the MMD context, Bodenham and Kawahara [5] leveraged the semigroup property of the Laplace kernel to achieve linear-time acceleration of the MMD mean (euMMD) for univariate settings. However, this computational advantage has not yet been extended to unbiased variance estimation, where handling complex second-order terms, including Frobenius norms of kernel matrices, norms of row-sum vectors, and cross-matrix interactions, prevents existing linear-recursion schemes from providing exact results [28].
To bridge this gap, we exploit the analytical properties of the Laplace kernel [16, 5] to reformulate all terms in the variance expression as a system of linear prefix sums. This reformulation enables exact computation of the unbiased MMD variance with quasi-linear time complexity $\mathcal{O}(n\log n)$.
The remainder of this paper is organized as follows. Section 2 reviews the necessary background, including the definition of MMD, U-statistics, and the Hoeffding decomposition, as well as existing unbiased MMD acceleration techniques based on sorting methods. Section 3 derives the variance expression via the Hoeffding decomposition and presents the corresponding unbiased estimator. Section 4 introduces an accelerated variance computation method based on the Laplace kernel and the sorting methods. Finally, Section 5 presents numerical experiments demonstrating the proposed method’s performance.
2 Background
2.1 Maximum Mean Discrepancy
Assume $\boldsymbol{X},\boldsymbol{X^{\prime}}{\sim}\mathbb{P}$ and $\boldsymbol{Y},\boldsymbol{Y^{\prime}}{\sim}\mathbb{Q}$ are independent copies. The maximum mean discrepancy is defined as [11, 22]:
$\mathrm{MMD}^{2}\left(\boldsymbol{X},\boldsymbol{Y}\right)=\mathbb{E}_{\boldsymbol{X},\boldsymbol{X^{\prime}}}\left[k\left(\boldsymbol{X},\boldsymbol{X^{\prime}}\right)\right]+\mathbb{E}_{\boldsymbol{Y},\boldsymbol{Y^{\prime}}}\left[k\left(\boldsymbol{Y},\boldsymbol{Y^{\prime}}\right)\right]-2\mathbb{E}_{\boldsymbol{X},\boldsymbol{Y}}\left[k\left(\boldsymbol{X},\boldsymbol{Y}\right)\right]$ (1)
Given i.i.d. finite samples $\mathbf{X}=\left\{x_{i}\right\}_{i=1}^{n}$ and $\mathbf{Y}=\left\{y_{j}\right\}_{j=1}^{m}$ from $\mathbb{P}$ and $\mathbb{Q}$, respectively, an unbiased estimator of the squared MMD is given by [6]:
$\widehat{\mathrm{MMD}}^{2}(\mathbf{X},\mathbf{Y})=\frac{1}{n(n-1)}\sum_{i\neq i^{\prime}}k(x_{i},x_{i^{\prime}})+\frac{1}{m(m-1)}\sum_{j\neq j^{\prime}}k(y_{j},y_{j^{\prime}})-\frac{2}{nm}\sum_{i,j}k(x_{i},y_{j})$ (2)
which is a linear combination of two single-sample U-statistics and one generalized U-statistic [26].
The effectiveness of MMD hinges on the use of characteristic kernels, for which the kernel mean embedding uniquely represents probability distributions in the associated reproducing kernel Hilbert space (RKHS) [27, 9]. Under this condition, MMD induces a proper metric on distributions, ensuring that any difference between them is detectable [8, 10].
In practice, translation-invariant kernels are particularly appealing due to their simple analytic form and favorable computational properties, as characterized by Bochner’s theorem. Two commonly used examples are the Gaussian kernel and the Laplace kernel
$k_{\mathrm{G}}(x,x^{\prime};\sigma)=\exp\left(-(x-x^{\prime})^{2}/{2\sigma^{2}}\right)$ (3)
$k_{\mathrm{L}}(x,x^{\prime};\sigma)=\exp\left(-{{|x-x^{\prime}|}/{\sigma}}\right)$ (4)
where the kernel parameter $\sigma$ controls the scale of the feature space and directly influences the test power. A common choice is the median heuristic [12], where $\sigma$ is set to the median of all pairwise sample distances. This approach requires no prior knowledge of the underlying distributions and is widely regarded as a standard baseline in the literature.
Finally, although $\widehat{\mathrm{MMD}}^{2}$ enjoys favorable theoretical properties, its computation relies on the explicit construction of kernel matrices, resulting in a computational complexity of $\mathcal{O}(d(n+m)^{2})$. Letting $N=n+m$, this complexity can be written as $\mathcal{O}(dN^{2})$ [5]. In large-scale datasets or scenarios where moderate sample sizes require repeated evaluations, such quadratic computational cost severely limits the practical applicability of MMD-based statistical inference.
2.2 U-statistics and Hoeffding Decomposition
Depending on the sample source, U-statistics can be classified into single-sample U-statistics and generalized U-statistics. Both types form the theoretical foundation of the MMD estimator [26]. Let finite samples be $\mathbf{X}=\{x_{i}\}_{i=1}^{n}$ i.i.d. from $\boldsymbol{X}\sim\mathbb{P}$ and $\mathbf{Y}=\{y_{j}\}_{j=1}^{m}$ i.i.d. from $\boldsymbol{Y}\sim\mathbb{Q}$, where $\boldsymbol{X}$ and $\boldsymbol{Y}$ are independent. Consider a function $\theta(\boldsymbol{X})=\mathbb{E}_{\boldsymbol{X}}[h(x_{1},\dots,x_{r})]$, whose unbiased estimator is
$U_{n}=\binom{n}{r}^{-1}\sum_{1\leq i_{1}<\dots<i_{r}\leq n}h(x_{i_{1}},\dots,x_{i_{r}}),$ (5)
where $r$ denotes the order of the U-statistic and $n$ the sample size. For the single-sample terms in the $\widehat{\mathrm{MMD}}^{2}$ estimator (Eq. 2), the kernel $k(x_{i},x_{i^{\prime}})$ acts on pairs of samples, corresponding to the case where $r=2$.
Define the conditional expectation projections as
$h_{c}(x_{i_{1}},\dots,x_{i_{c}})=\int h(x_{i_{1}},\dots,x_{i_{c}},x_{i_{c+1}},\dots,x_{i_{r}})\,d\mathbb{P}(x_{i_{c+1}})\cdots d\mathbb{P}(x_{i_{r}}),\quad 1\leq c\leq r.$ (6)
and $\tilde{h}(x_{i_{1}},\dots,x_{i_{r}})=h(x_{i_{1}},\dots,x_{i_{r}})-\theta(\boldsymbol{X})$ is the centered kernel.
The orthogonal projection components $g_{c}$ are recursively defined by [14, 4]:
$g_{1}(x_{i_{1}})$ $=\tilde{h}_{1}(x_{i_{1}}),$
$g_{2}(x_{i_{1}},x_{i_{2}})$ $=\tilde{h}_{2}(x_{i_{1}},x_{i_{2}})-g_{1}(x_{i_{1}})-g_{1}(x_{i_{2}}),$
$g_{c}(x_{i_{1}},\dots,x_{i_{c}})$ $=\tilde{h}_{c}(x_{i_{1}},\dots,x_{i_{c}})-\sum_{s=1}^{c-1}\sum_{1\leq i_{1}<\dots<i_{s}\leq c}g_{s}(x_{i_{1}},\dots,x_{i_{s}}),\quad c\geq 3.$ (7)
These components satisfy the zero-mean property: $\mathbb{E}_{\boldsymbol{X}}[g_{c}(x_{i_{1}},\dots,x_{i_{c}})]=0$.
Define $S_{c,n}=\sum_{1\leq i_{1}<\cdots<i_{c}\leq n}g_{c}(x_{i_{1}},\dots,x_{i_{c}})$, then the Hoeffding decomposition yields by [15]:
$U_{n}-\theta(\boldsymbol{X})=\sum_{c=1}^{r}\binom{r}{c}\binom{n}{c}^{-1}S_{c,n}.$ (8)
For parameters involving interactions between two distributions $\theta(\boldsymbol{X},\boldsymbol{Y})=\mathbb{E}[h(x_{1},\dots,x_{r_{1}};y_{1},\dots,y_{r_{2}})]$, the corresponding generalized U-statistic is defined as
$U_{n,m}=\frac{1}{\binom{n}{r_{1}}\binom{m}{r_{2}}}\sum_{1\leq i_{1}<\cdots<i_{r_{1}}\leq n}\sum_{1\leq j_{1}<\cdots<j_{r_{2}}\leq m}h(x_{i_{1}},\dots,x_{i_{r_{1}}};y_{j_{1}},\dots,y_{j_{r_{2}}}),$ (9)
where $r_{1}$ and $r_{2}$ denote the orders of the U-statistic for the two samples, respectively. For the cross-sample interaction term in the $\widehat{\mathrm{MMD}}^{2}$ estimator, the kernel $k(x_{i},y_{j})$ involves a single observation from each distribution, which corresponds to the case where $r_{1}=1$ and $r_{2}=1$.
For the generalized U-statistic $U_{n,m}$ defined in Eq. 9, its asymptotic variance are discussed in [20]. Following the recursive logic of the single-sample case, we specify the orthogonal projection components $g_{c_{1},c_{2}}$ for the generalized case as follows.
Let $\tilde{h}=h-\theta(\boldsymbol{X},\boldsymbol{Y})$ be the centered kernel. For $0\leq c_{1}\leq r_{1}$ and $0\leq c_{2}\leq r_{2}$, we define the joint conditional expectation projections as:
$\tilde{h}_{c_{1},c_{2}}(x_{i_{1}},\dots,x_{i_{c_{1}}};y_{j_{1}},\dots,y_{j_{c_{2}}})=\mathbb{E}[\tilde{h}\mid x_{i_{1}},\dots,x_{i_{c_{1}}},y_{j_{1}},\dots,y_{j_{c_{2}}}].$ (10)
The orthogonal projection components $g_{c_{1},c_{2}}$ are recursively defined as:
$g_{c_{1},c_{2}}(x_{i_{1}},\dots,x_{i_{c_{1}}};y_{j_{1}},\dots,y_{j_{c_{2}}})=\tilde{h}_{c_{1},c_{2}}-\sum_{\begin{subarray}{c}0\leq s_{1}\leq c_{1}\\ 0\leq s_{2}\leq c_{2}\\ 0<s_{1}+s_{2}<c_{1}+c_{2}\end{subarray}}\sum_{\mathscr{S}_{s_{1},s_{2}}}g_{s_{1},s_{2}}(\cdot),$ (11)
where $\mathscr{S}_{s_{1},s_{2}}$ represents all subsets of size $s_{1}$ from $\{x_{i_{1}},\dots,x_{i_{c_{1}}}\}$ and size $s_{2}$ from $\{y_{j_{1}},\dots,y_{j_{c_{2}}}\}$. These components satisfy $\mathbb{E}[g_{c_{1},c_{2}}]=0$ for all $c_{1}+c_{2}>0$.
Define the partial sums as $S_{c_{1},c_{2},n,m}=\sum_{1\leq i_{1}<\dots<i_{c_{1}}\leq n}\sum_{1\leq j_{1}<\dots<j_{c_{2}}\leq m}g_{c_{1},c_{2}}(x_{i_{1}},\dots,x_{i_{c_{1}}};y_{j_{1}},\dots,y_{j_{c_{2}}})$, the generalized Hoeffding decomposition is then expressed as:
$U_{n,m}-\theta(\boldsymbol{X},\boldsymbol{Y})=\sum_{\begin{subarray}{c}c_{1}=0,\dots,r_{1}\\ c_{2}=0,\dots,r_{2}\\ (c_{1},c_{2})\neq(0,0)\end{subarray}}\binom{r_{1}}{c_{1}}\binom{r_{2}}{c_{2}}\binom{n}{c_{1}}^{-1}\binom{m}{c_{2}}^{-1}S_{c_{1},c_{2},n,m}.$ (12)
2.3 Lossless acceleration of the unbiased MMD estimator with the Laplace kernel
For univariate samples, directly computing Eq. 2 requires $\mathcal{O}(N^{2})$ computational complexity. Based on several lemmas for the Laplacian kernel presented in Appendix A, Bodenham and Kawahara [5] proposed a lossless acceleration strategy that exploits the ordering structure of the samples, reducing the complexity to $\mathcal{O}(N\log N)$. The procedure is summarized as follows.
First, by applying Lemmas 1 and 2 in Appendix A, Eq. 2 can be rewritten in the following form:
$\widehat{\mathrm{MMD}}_{\mathrm{eu}}^{2}\left(\mathbf{X},\mathbf{Y}\right)=2\left(\alpha_{1}T_{1}+\alpha_{2}T_{2}-\alpha_{3}T_{3}\right)$ (13)
where the coefficients are defined as $\alpha_{1}={1}\big/[{n\left(n-1\right)}]$, $\alpha_{2}={1}\big/[{m\left(m-1\right)}]$, $\alpha_{3}={1}\big/[{nm}]$. The summation terms correspond to the lower triangular parts of the kernel matrices and are given by $T_{1}=\sum_{i=2}^{n}{\sum_{i^{\prime}=1}^{i-1}{k\left(x_{i},x_{i^{\prime}}\right)}}$, $T_{2}=\sum_{j=2}^{m}{\sum_{j^{\prime}=1}^{j-1}{k\left(y_{j},y_{j^{\prime}}\right)}}$, $T_{3}=\sum_{i=1}^{n}{\sum_{j=1}^{m}{k\left(x_{i},y_{j}\right)}}$.
To compute $T_{1}$, the sample sequence $\mathbf{X}$ is first sorted in ascending order, i.e., $x_{1}\leq x_{2}\leq\cdots\leq x_{n}$. Define the forward cumulative sum $R_{i}^{\mathbf{X}}=\sum_{i^{\prime}=1}^{i-1}{k\left(x_{i},x_{i^{\prime}}\right)}$ and the exponential difference between adjacent samples $D_{i}^{\mathbf{X}}=\exp\left(-{{\left(x_{i}-x_{i-1}\right)}/{\sigma}}\right)$. According to Proposition 4 in the appendix A, the following recurrence relation holds:
$R_{i}^{\mathbf{X}}=\left(R_{i-1}^{\mathbf{X}}+1\right)\cdot D_{i}^{\mathbf{X}}.$ (14)
As a result, $T_{1}$ can be computed in linear time as
$T_{1}=\sum_{i=2}^{n}{R_{i}^{\mathbf{X}}}.$ (15)
Similarly, for $T_{2}$, sorting $\mathbf{Y}$ in ascending order $y_{1}\leq y_{2}\leq\cdots\leq y_{m}$ yields
$T_{2}=\sum_{j=2}^{m}{R_{i}^{\mathbf{Y}}}.$ (16)
For the cross-sample term $T_{3}$, a merged-set strategy is adopted. Let $\mathbf{Z}=\mathbf{X}\cup\mathbf{Y}$ be the combined sample set, and sort it as $z_{1}\leq z_{2}\leq\cdots\leq z_{n+m}$. Define the total lower-triangular sum over the merged set as $T_{4}=\sum_{k=2}^{n+m}{\sum_{k^{\prime}=1}^{k-1}{k\left(z_{k},z_{k^{\prime}}\right)}}$. By set decomposition, $T_{4}$ contains all terms in $T_{1}$, $T_{2}$, and $T_{3}$, and thus $T_{3}=T_{4}-T_{1}-T_{2}$.
The quantity $T_{4}$ can be efficiently computed using the same linear recurrence $R_{k}^{\boldsymbol{Z}}=\left(R_{k-1}^{\boldsymbol{Z}}+1\right)\cdot D_{k}^{\boldsymbol{Z}}$, which gives
$T_{4}=\sum_{k=2}^{n+m}{R_{k}^{\boldsymbol{Z}}}.$ (17)
The core logic of the above procedure is summarized in Algorithm 1, which achieves a computational complexity of $\mathcal{O}(n\log n)$.
Algorithm 1 TriSSL: Triangular Sorted Sum of Laplacians
1:Ordered sample sequence $\mathbf{S}=\{s_{1},\dots,s_{l}\}$, kernel parameter $\sigma$
2:Lower-triangular summation result $T$
[…]
4:for $i=2$ to $l$ do
5:  $D_{i}\Leftarrow\exp\bigl(-(s_{i}-s_{i-1})/\sigma\bigr)$
6:  $R_{i}\Leftarrow(R_{i-1}+1)\cdot D_{i}$
7:  $T\Leftarrow T+R_{i}$
8:end for
9:return $T$
3 Unbiased estimator of Maximum Mean Discrepancy variance
3.1 Variance derivation of the unbiased MMD estimator
Based on the theory of U-statistics and the Hoeffding decomposition presented in Section 2.2, this subsection provides a rigorous derivation of the variance of the unbiased estimator $\widehat{\mathrm{MMD}}^{2}$.
First, we define the components in Eq. 1 as
$\theta_{\boldsymbol{X}}:=\mathbb{E}_{\boldsymbol{X},\boldsymbol{X^{\prime}}}\left[k\left(\boldsymbol{X},\boldsymbol{X^{\prime}}\right)\right];\,\theta_{\boldsymbol{Y}}:=\mathbb{E}_{\boldsymbol{Y},\boldsymbol{Y^{\prime}}}\left[k\left(\boldsymbol{Y},\boldsymbol{Y^{\prime}}\right)\right];\,\theta_{\boldsymbol{XY}}:=\mathbb{E}_{\boldsymbol{X},\boldsymbol{Y}}\left[k\left(\boldsymbol{X},\boldsymbol{Y}\right)\right].$
Correspondingly, the components in Eq. 2 are denoted as
$A:=\frac{1}{n(n-1)}\sum_{i\neq i^{\prime}}k(x_{i},x_{i^{\prime}});\,B:=\frac{1}{m(m-1)}\sum_{j\neq j^{\prime}}k(y_{j},y_{j^{\prime}});\,C:=\frac{1}{nm}\sum_{i,j}k(x_{i},y_{j}).$
According to the Hoeffding decomposition theorem (Eq. 8 and 12), the deviation of each component can be expanded into a sum of linear terms and second-order residual terms.
For the within-$\boldsymbol{X}$ component,
$A-\theta_{\boldsymbol{X}}=\frac{2}{n}\sum_{i=1}^{n}{g_{1,A}\left(x_{i}\right)}+\frac{2}{n\left(n-1\right)}\sum_{1\leq i<i^{\prime}\leq n}{g_{2,A}\left(x_{i},x_{i^{\prime}}\right)}$ (18)
where $g_{1,A}\left(x_{i}\right)=\mathbb{E}_{\boldsymbol{X^{\prime}}}\left[k\left(x_{i},\boldsymbol{X^{\prime}}\right)\right]-\theta_{\boldsymbol{X}}$, $g_{2,A}\left(x_{i},x_{i^{\prime}}\right)=k\left(x_{i},x_{i^{\prime}}\right)-g_{1,A}\left(x_{i}\right)-g_{1,A}\left(x_{i^{\prime}}\right)-\theta_{\boldsymbol{X}}$.
Similarly, for the within-$\boldsymbol{Y}$ component,
$B-\theta_{\boldsymbol{Y}}=\frac{2}{m}\sum_{j=1}^{m}{g_{1,B}\left(y_{j}\right)}+\frac{2}{m\left(m-1\right)}\sum_{1\leq j<j^{\prime}\leq m}{g_{2,B}\left(y_{j},y_{j^{\prime}}\right)}$ (19)
where $g_{1,B}\left(y_{j}\right)=\mathbb{E}_{\boldsymbol{Y}^{\prime}}\left[k\left(y_{j},\boldsymbol{Y}^{\prime}\right)\right]-\theta_{\boldsymbol{Y}}$, $g_{2,B}\left(y_{j},y_{j^{\prime}}\right)=k\left(y_{j},y_{j^{\prime}}\right)-g_{1,B}\left(y_{j}\right)-g_{1,B}\left(y_{j^{\prime}}\right)-\theta_{\boldsymbol{Y}}$.
For the cross-sample component,
$C-\theta_{\boldsymbol{XY}}=\frac{1}{n}\sum_{i=1}^{n}{g_{1,CX}\left(x_{i}\right)}+\frac{1}{m}\sum_{j=1}^{m}{g_{1,CY}\left(y_{j}\right)}+\frac{1}{nm}\sum_{i=1}^{n}{\sum_{j=1}^{m}{g_{2,C}\left(x_{i},y_{j}\right)}}$ (20)
where $g_{1,CX}\left(x_{i}\right)=\mathbb{E}_{\boldsymbol{Y}}\left(k\left(x_{i},\boldsymbol{Y}\right)\right)-\theta_{\boldsymbol{XY}}$, $g_{1,CY}\left(y_{j}\right)=\mathbb{E}_{\boldsymbol{X}}\left(k\left(\boldsymbol{X},y_{j}\right)\right)-\theta_{\boldsymbol{XY}}$, $g_{2,C}\left(x_{i},y_{j}\right)=k\left(x_{i},y_{j}\right)-g_{1,CX}\left(x_{i}\right)-g_{1,CY}\left(y_{j}\right)-\theta_{\boldsymbol{XY}}$.
Unlike the approach in [30], which directly applies the general properties of generalized U-statistics by treating the entire MMD as a single monolithic entity, we leverage the natural form of the MMD estimator as a combination of distinct U-statistic components. By combining Eq. 18, 19, and 20, the deviation of the unbiased estimator can be expressed as:
$\widehat{\mathrm{MMD}}^{2}-\mathrm{MMD}^{2}=\mathcal{T}_{1}+\mathcal{T}_{2}$ (21)
here, $\mathcal{T}_{1}$ denotes the sum of first-order projection terms,
$\mathcal{T}_{1}=\frac{2}{n}\sum_{i=1}^{n}{\left(\mathcal{U}\left(x_{i}\right)-\mathbb{E}\left[\mathcal{U}\left(x_{i}\right)\right]\right)}+\frac{2}{m}\sum_{j=1}^{m}{\left(\mathcal{V}\left(y_{j}\right)-\mathbb{E}\left[\mathcal{V}\left(y_{j}\right)\right]\right)}$ (22)
where the influence functions are defined as
$\mathcal{U}\left(x_{i}\right)$ $:=\mathbb{E}_{\boldsymbol{X}^{\prime}}\left[k\left(x_{i},\boldsymbol{X}^{\prime}\right)\right]-\mathbb{E}_{\boldsymbol{Y}}\left[k\left(x_{i},\boldsymbol{Y}\right)\right]$ (23)
$\mathcal{V}\left(y_{j}\right)$ $:=\mathbb{E}_{\boldsymbol{Y}^{\prime}}\left[k\left(y_{j},\boldsymbol{Y}^{\prime}\right)\right]-\mathbb{E}_{\boldsymbol{X}}\left[k\left(\boldsymbol{X},y_{j}\right)\right]$ (24)
The term $\mathcal{T}_{2}$ corresponds to the second-order residual terms,
$\mathcal{T}_{2}=\frac{2}{n\left(n-1\right)}\zeta_{2,A}+\frac{2}{m\left(m-1\right)}\zeta_{2,B}-\frac{2}{nm}\zeta_{2,C}$ (25)
where $\zeta_{2,A}=\sum_{1\leq i<i^{\prime}\leq n}{g_{2,A}\left(x_{i},x_{i^{\prime}}\right)}$, $\zeta_{2,B}=\sum_{1\leq j<j^{\prime}\leq m}{g_{2,B}\left(y_{j},y_{j^{\prime}}\right)}$, $\zeta_{2,C}=\sum_{i=1}^{n}{\sum_{j=1}^{m}{g_{2,C}\left(x_{i},y_{j}\right)}}$.
Since different orders in the Hoeffding decomposition are orthogonal, i.e., $\mathrm{Cov}(H_{1},H_{2})=0$ [15], the variance of $\widehat{\mathrm{MMD}}^{2}$ admits the exact decomposition
$\operatorname{Var}\left(\widehat{\mathrm{MMD}}^{2}\right)=\mathrm{Var}\left(\mathcal{T}_{1}\right)+\mathrm{Var}\left(\mathcal{T}_{2}\right)$ (26)
where the variance of the linear term $\operatorname{Var}(\mathcal{T}_{1})$ captures the dominant fluctuations of the sample means, while the variance of the remainder term $\operatorname{Var}(\mathcal{T}_{2})$ characterizes the second-order interactions induced by the kernel.
3.2 Simplification and unbiased estimation of the variance
3.2.1 Simplification and unbiased estimation of first-order projection terms
To translate the theoretical decomposition in Eq. 26 into a practical and computable form, we leverage the fundamental independence between the two sample sets, and the orthogonal representation allows for a decoupled estimation of the variance components. Specifically, for the first-order term $\mathcal{T}_{1}$, the mutual independence of $\boldsymbol{X}$ and $\boldsymbol{Y}$ ensures that the covariance between the sample-specific fluctuations vanishes. Consequently, the variance of the linear term simplifies to the weighted sum of marginal variances:
$\widehat{\operatorname{Var}}(\mathcal{T}_{1})=\frac{4(n-2)}{n(n-1)}\widehat{\operatorname{Var}}(\mathcal{U})+\frac{4(m-2)}{m(m-1)}\widehat{\operatorname{Var}}(\mathcal{V}),$ (27)
where $\widehat{\operatorname{Var}}(\mathcal{U})$ and $\widehat{\operatorname{Var}}(\mathcal{V})$ represent the variances of population projections. To realize this empirically, we replace these population expectations with their empirical projections:
For a sample point $x_{i}\in\mathbf{X}$, the continuous projection $\mathcal{U}(x_{i})$ is discretized using sample averages as
$\hat{\mathcal{U}}(x_{i})=\frac{1}{n-1}\sum_{i^{\prime}\neq i}k(x_{i},x_{i^{\prime}})-\frac{1}{m}\sum_{j=1}^{m}k(x_{i},y_{j})$ (28)
The unbiased estimator of the variance $\operatorname{Var}(\mathcal{U}(X))$ is therefore given by the sample variance:
$\widehat{\operatorname{Var}}(\mathcal{U})=\frac{1}{n-1}\sum_{i=1}^{n}\left(\hat{\mathcal{U}}(x_{i})-\bar{\hat{\mathcal{U}}}\right)^{2}$ (29)
where $\bar{\hat{\mathcal{U}}}=\sum_{i=1}^{n}\hat{\mathcal{U}}(x_{i})\big/n$ and an analogous estimator can be obtained for $\hat{\mathcal{V}}(y_{j})$ and $\widehat{\operatorname{Var}}(\mathcal{V})$:
$\hat{\mathcal{V}}(y_{j})=\frac{1}{m-1}\sum_{j^{\prime}\neq j}k(y_{j},y_{j^{\prime}})-\frac{1}{n}\sum_{i=1}^{n}k(x_{i},y_{j})$ (30)
$\widehat{\operatorname{Var}}(\mathcal{V})=\frac{1}{m-1}\sum_{j=1}^{m}\left(\hat{\mathcal{V}}(y_{j})-\bar{\hat{\mathcal{V}}}\right)^{2}$ (31)
where we avoid the appearance of complex cross-covariance terms that are prone to estimation bias in finite samples (as seen in [30]). By projecting the estimator onto orthogonal Hoeffding components, we decompose the variance into a streamlined structure of marginal variances and interaction residuals. This not only simplifies the analytical form but also ensures the robustness of the estimator in imbalanced scenarios where $n\neq m$.
3.2.2 Simplification and unbiased estimation of second-order residual terms
Parallel to the treatment of the linear term, the variance of the second-order remainder $\mathcal{T}_{2}$ can be simplified by exploiting the mutual orthogonality of the $g_{2}$ components across different sample pairs. Since $g_{2,A}$, $g_{2,B}$, and $g_{2,C}$ are constructed to be zero-mean and mutually uncorrelated under the independent sampling of $\boldsymbol{X}$ and $\boldsymbol{Y}$, the variance $\operatorname{Var}(\mathcal{T}_{2})$ avoids any cross-product terms. This allows us to expand the second-order variance into a weighted sum of the dispersion of each interaction kernel:
$\operatorname{Var}(\mathcal{T}_{2})=\operatorname{Var}\left(\frac{2\zeta_{2,A}}{n(n-1)}\right)+\operatorname{Var}\left(\frac{2\zeta_{2,B}}{m(m-1)}\right)+\operatorname{Var}\left(\frac{2\zeta_{2,C}}{nm}\right)$ (32)
As an illustrative example for the within-sample component $A$, consider the second-order sum $\zeta_{2,A}=\sum_{1\leq i<i^{\prime}\leq n}g_{2,A}(x_{i},x_{i^{\prime}})$. Since $g_{2,A}$ is a centered orthogonal projection component, for any two distinct index pairs $i,i^{\prime}$} and $j,j^{\prime}$}, if the pairs do not coincide exactly, then $\mathrm{Cov}(g_{2,A}(x_{i},x_{i^{\prime}}),g_{2,A}(x_{j},x_{j^{\prime}}))=0$. As a consequence, the variance of the sum equals the sum of the variances of the individual terms:
$\operatorname{Var}\left(\frac{2\zeta_{2,A}}{n(n-1)}\right)$ $=\frac{4}{n^{2}(n-1)^{2}}\cdot\binom{n}{2}\operatorname{Var}(g_{2,A})$
$=\frac{4}{n^{2}(n-1)^{2}}\cdot\frac{n(n-1)}{2}\mathbb{E}[g_{2,A}^{2}]=\frac{2}{n(n-1)}\mathbb{E}[g_{2,A}^{2}]$ (33)
Similarly, for the cross-sample mixed component $C$, since $x_{i}$ and $y_{j}$ are independent, the covariance is non-zero only when the indices coincide exactly:
$\operatorname{Var}\left(\frac{2\zeta_{2,C}}{nm}\right)=\frac{4}{n^{2}m^{2}}\cdot(nm)\mathbb{E}[g_{2,C}^{2}]=\frac{4}{nm}\mathbb{E}[g_{2,C}^{2}]$ (34)
At this stage, Eq. 32 can be reformulated as
$\operatorname{Var}(\mathcal{T}_{2})=\frac{2}{n(n-1)}\mathbb{E}[g_{2,A}^{2}]+\frac{2}{m(m-1)}\mathbb{E}[g_{2,B}^{2}]+\frac{4}{nm}\mathbb{E}[g_{2,C}^{2}]$ (35)
To further simplify, Consider the orthogonality of first-order components (Section 2.2) and the nesting property of conditional expectations, $\mathbb{E}[\boldsymbol{Z}\cdot\mathbb{E}[\boldsymbol{Z}|\boldsymbol{X}]]=\mathbb{E}[(\mathbb{E}[\boldsymbol{Z}|\boldsymbol{X}])^{2}]$, which allows the squared second-order orthogonal components to be simplified. Taking $g_{2,A}$ as an example, where $g_{2,A}(x,x^{\prime})=k(x,x^{\prime})-g_{1,A}(x)-g_{1,A}(x^{\prime})-\theta_{\boldsymbol{X}}$. The expectation of the squared term and expanding yields
$\mathbb{E}[g_{2,A}^{2}]$ $=\mathbb{E}[k(x,x^{\prime})^{2}]+\mathbb{E}[g_{1,A}(x)^{2}]+\mathbb{E}[g_{1,A}(x^{\prime})^{2}]+\theta_{\boldsymbol{X}}^{2}$ (36)
$-2\mathbb{E}[k(x,x^{\prime})\cdot g_{1,A}(x)]-2\mathbb{E}[k(x,x^{\prime})\cdot g_{1,A}(x^{\prime})]-2\theta_{\boldsymbol{X}}\mathbb{E}[k(x,x^{\prime})]$ (37)
$+2\mathbb{E}[g_{1,A}(x)g_{1,A}(x^{\prime})]-2\theta_{\boldsymbol{X}}\mathbb{E}[g_{1,A}(x)]-2\theta_{\boldsymbol{X}}\mathbb{E}[g_{1,A}(x^{\prime})]$ (38)
Using the facts that $\mathbb{E}[g_{1,A}(x)]=0$, $\mathbb{E}[k(x,x^{\prime})\cdot g_{1,A}(x)]=\mathbb{E}[g_{1,A}(x)^{2}]$, and $\mathbb{E}[g_{1,A}(x)g_{1,A}(x^{\prime})]=0$, the above expression simplifies to
$\mathbb{E}[g_{2,A}^{2}]=\mathbb{E}[k(x,x^{\prime})^{2}]-2\mathbb{E}_{\boldsymbol{X}}\left[\left(\mathbb{E}_{\boldsymbol{X^{\prime}}}[k(\boldsymbol{X},\boldsymbol{X^{\prime}})]\right)^{2}\right]+\theta_{\boldsymbol{X}}^{2},$ (39)
analogous derivation can be obtained for $\mathbb{E}[g_{2,B}^{2}]$:
$\mathbb{E}[g_{2,B}^{2}]=\mathbb{E}[k(y,y^{\prime})^{2}]-2\mathbb{E}_{\boldsymbol{Y}}\left[\left(\mathbb{E}_{\boldsymbol{Y^{\prime}}}[k(\boldsymbol{Y},\boldsymbol{Y^{\prime}})]\right)^{2}\right]+\theta_{\boldsymbol{Y}}^{2}$ (40)
Consider the mixed component $g_{2,C}(x,y)=k(x,y)-g_{1,CX}(x)-g_{1,CY}(y)-\theta_{\boldsymbol{XY}}$, which arises from the interaction between samples $\mathbf{X}$ and $\mathbf{Y}$. Since $\mathbb{P}$ and $\mathbb{Q}$ are independent, the first-order projection terms $g_{1,CX}(x)$ and $g_{1,CY}(y)$ are also independent. Expanding the squared term and eliminating zero-mean cross terms yields
$\mathbb{E}[g_{2,C}^{2}]$ $=\mathbb{E}[k(x,y)^{2}]+\mathbb{E}[g_{1,CX}^{2}]+\mathbb{E}[g_{1,CY}^{2}]+\theta_{\boldsymbol{XY}}^{2}$
$-2\mathbb{E}[k(x,y)g_{1,CX}(x)]-2\mathbb{E}[k(x,y)g_{1,CY}(y)]+2\theta_{\boldsymbol{XY}}\mathbb{E}[k(x,y)]$ (41)
Using again the projection property $\mathbb{E}[k(x,y)g_{1,CX}(x)]=\mathbb{E}[g_{1,CX}^{2}+\theta_{\boldsymbol{XY}}g_{1,CX}]=\mathbb{E}[g_{1,CX}^{2}]$, the expression simplifies to
$\mathbb{E}[g_{2,C}^{2}]=\mathbb{E}[k(x,y)^{2}]-\mathbb{E}_{\boldsymbol{X}}\left[\left(\mathbb{E}_{\boldsymbol{Y}}[k(\boldsymbol{X},\boldsymbol{Y})]\right)^{2}\right]-\mathbb{E}_{\boldsymbol{Y}}\left[\left(\mathbb{E}_{\boldsymbol{X}}[k(\boldsymbol{X},\boldsymbol{Y})]\right)^{2}\right]+\theta_{\boldsymbol{XY}}^{2}$ (42)
By applying the inclusion-exclusion principle, the unbiased estimator can be expressed entirely in terms of unconstrained matrix operations [28]. Specifically, using the falling factorial $(n)_{k}$ and the centered kernel matrices $\tilde{\mathbf{K}}_{XX}$ and $\tilde{\mathbf{K}}_{YY}$ (with zero diagonals), the expectation $\mathbb{E}[g_{2}^{2}]$ can be written as a linear combination of the squared Frobenius norms, the squared row and column sums, and the squared total sum of the kernel matrices.
$\widehat{\mathbb{E}[g_{2,A}^{2}]}$ $=\frac{1}{(n)_{2}}\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}-\frac{2}{(n)_{3}}\left(\|\tilde{\mathbf{K}}_{XX}\mathbf{1}\|_{2}^{2}-\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}\right)$
$+\frac{1}{(n)_{4}}\left((\mathbf{1}^{\top}\tilde{\mathbf{K}}_{XX}\mathbf{1})^{2}-4\|\tilde{\mathbf{K}}_{XX}\mathbf{1}\|_{2}^{2}+2\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}\right)$ (43)
$\widehat{\mathbb{E}[g_{2,B}^{2}]}$ $=\frac{1}{(m)_{2}}\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}-\frac{2}{(m)_{3}}\left(\|\tilde{\mathbf{K}}_{YY}\mathbf{1}\|_{2}^{2}-\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}\right)$
$+\frac{1}{(m)_{4}}\left((\mathbf{1}^{\top}\tilde{\mathbf{K}}_{YY}\mathbf{1})^{2}-4\|\tilde{\mathbf{K}}_{YY}\mathbf{1}\|_{2}^{2}+2\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}\right)$ (44)
$\widehat{\mathbb{E}[g_{2,C}^{2}]}$ $=\frac{1}{nm}\|\mathbf{K}_{XY}\|_{F}^{2}-\frac{1}{n(m)_{2}}\left(\|\mathbf{K}_{XY}^{\top}\mathbf{1}_{n}\|_{2}^{2}-\|\mathbf{K}_{XY}\|_{F}^{2}\right)-\frac{1}{m(n)_{2}}\left(\|\mathbf{K}_{XY}\mathbf{1}_{m}\|_{2}^{2}-\|\mathbf{K}_{XY}\|_{F}^{2}\right)$
$+\frac{1}{(n)_{2}(m)_{2}}\left((\mathbf{1}_{n}^{\top}\mathbf{K}_{XY}\mathbf{1}_{m})^{2}-\|\mathbf{K}_{XY}\mathbf{1}_{m}\|_{2}^{2}-\|\mathbf{K}_{XY}^{\top}\mathbf{1}_{n}\|_{2}^{2}+\|\mathbf{K}_{XY}\|_{F}^{2}\right)$ (45)
where $(n)_{k}={n!}\big/{(n-k)!}$ denotes the falling factorial.
3.2.3 The unbiased estimator of MMD variance
Using the simplification results provided in Eq. 29314344, and 45, the full analytical expression of $\operatorname{Var}(\widehat{\mathrm{MMD}}^{2})$ is given by
$\widehat{\operatorname{Var}}\left(\widehat{\mathrm{MMD}}^{2}\right)$ $=\frac{4(n-2)}{n(n-1)}\widehat{\operatorname{Var}}(\mathcal{U})+\frac{4(m-2)}{m(m-1)}\widehat{\operatorname{Var}}(\mathcal{V})$
$+\frac{2}{n\left(n-1\right)}\widehat{\mathbb{E}[g_{2,A}^{2}]}+\frac{2}{m\left(m-1\right)}\widehat{\mathbb{E}[g_{2,B}^{2}]}+\frac{4}{nm}\widehat{\mathbb{E}[g_{2,C}^{2}]}$
$=\frac{4(n-2)}{n(n-1)^{2}}\sum_{i=1}^{n}\left(\hat{\mathcal{U}}(x_{i})-\bar{\hat{\mathcal{U}}}\right)^{2}+\frac{4(m-2)}{m(m-1)^{2}}\sum_{j=1}^{m}\left(\hat{\mathcal{V}}(y_{j})-\bar{\hat{\mathcal{V}}}\right)^{2}$
$+\frac{2}{(n)_{2}^{2}}\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}+\frac{2}{(m)_{2}^{2}}\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}+\frac{4}{n^{2}m^{2}}\|\mathbf{K}_{XY}\|_{F}^{2}$
$-\frac{4}{(n)_{2}(n)_{3}}\left(\|\tilde{\mathbf{K}}_{XX}\mathbf{1}\|_{2}^{2}-\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}\right)-\frac{4}{(m)_{2}(m)_{3}}\left(\|\tilde{\mathbf{K}}_{YY}\mathbf{1}\|_{2}^{2}-\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}\right)$
$-\frac{4}{n^{2}m(m)_{2}}\left(\|\mathbf{K}_{XY}^{\top}\mathbf{1}_{n}\|_{2}^{2}-\|\mathbf{K}_{XY}\|_{F}^{2}\right)-\frac{4}{m^{2}n(n)_{2}}\left(\|\mathbf{K}_{XY}\mathbf{1}_{m}\|_{2}^{2}-\|\mathbf{K}_{XY}\|_{F}^{2}\right)$
$+\frac{2}{(n)_{2}(n)_{4}}\left((\mathbf{1}^{\top}\tilde{\mathbf{K}}_{XX}\mathbf{1})^{2}-4\|\tilde{\mathbf{K}}_{XX}\mathbf{1}\|_{2}^{2}+2\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}\right)$
$+\frac{2}{(m)_{2}(m)_{4}}\left((\mathbf{1}^{\top}\tilde{\mathbf{K}}_{YY}\mathbf{1})^{2}-4\|\tilde{\mathbf{K}}_{YY}\mathbf{1}\|_{2}^{2}+2\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}\right)$
$+\frac{4}{n(n)_{2}m(m)_{2}}\left((\mathbf{1}_{n}^{\top}\mathbf{K}_{XY}\mathbf{1}_{m})^{2}-\|\mathbf{K}_{XY}\mathbf{1}_{m}\|_{2}^{2}-\|\mathbf{K}_{XY}^{\top}\mathbf{1}_{n}\|_{2}^{2}+\|\mathbf{K}_{XY}\|_{F}^{2}\right)$ (46)
Unlike previous works that rely on large-sample asymptotic approximations [30], our derived estimator is exactly unbiased for any finite sample sizes $n$ and $m$. By employing a combinatorial approach based on falling factorials $(n)_{k}$, we explicitly cancel out the systematic bias that typically arises when sample means are substituted for population expectations. This makes our estimator particularly robust in small-sample regimes or highly imbalanced settings ($n\ll m$).
Remark 1.
Under the null hypothesis $\mathcal{H}_{0}:\mathbb{P}=\mathbb{Q}$, the unbiased MMD statistic exhibits a "degenerate" behavior. In this regime, the population projections vanish, i.e.,
$\mathcal{U}\left(x\right)$ $:=\mathbb{E}_{\boldsymbol{X}^{\prime}}\left[k\left(x,\boldsymbol{X}^{\prime}\right)\right]-\mathbb{E}_{\boldsymbol{Y}}\left[k\left(x,\boldsymbol{Y}\right)\right]\equiv 0$ (47)
$\mathcal{V}\left(y\right)$ $:=\mathbb{E}_{\boldsymbol{Y}^{\prime}}\left[k\left(y,\boldsymbol{Y}^{\prime}\right)\right]-\mathbb{E}_{\boldsymbol{X}}\left[k\left(\boldsymbol{X},y\right)\right]\equiv 0,$ (48)
causing the $O(n^{-1})$ first-order variance $\operatorname{Var}(\mathcal{T}_{1})$ to drop to zero. Consequently, the total population variance is determined solely by the $O(n^{-2})$ second-order residual term $\mathcal{T}_{2}$ [11, 30].
Crucially, this theoretical degeneracy does not imply that the empirical variance of $\hat{\mathcal{U}}(x_{i})$ and $\hat{\mathcal{V}}(y_{j})$ disappears. In finite samples, the empirical projections (Eqs. 28 and 30) retain fluctuations due to the inherent coupling of sample pairs. Our estimator in Eq. (3.2.3) incorporates these fluctuations by integrating $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ into a single analytical form. As a result, our variance estimate is generally larger than those based on asymptotic approximations (e.g., [11]), as it explicitly captures high-order coupling terms that are typically neglected. This "unified" property is a significant advantage: under $\mathcal{H}_{1}$, the $O(n^{-1})$ terms naturally dominate, while under $\mathcal{H}_{0}$, the $O(n^{-2})$ terms automatically take over, which provides a more conservative and robust variance estimate.
4 Accelerated computation of Maximum Mean Discrepancy
4.1 Several Propositions and Lemma for acceleration
Under the Laplace kernel $k_{\mathrm{L}}(x,x^{\prime})=\exp(-{|x-x^{\prime}|}/{\sigma})$ and the Lemmas in the Appendix A, the following Propositions and Lemma hold.
Proposition 1.
Let $\{x_{1},\dots,x_{n}\}\subset\mathbb{R}$ be sorted such that $x_{1}\leq x_{2}\leq\dots\leq x_{n}$, and consider the Laplacian kernel. Define $R_{1}=L_{n}=0$. For $i=2,\dots,n$ and $i=n-1,\dots,1$, respectively, let
$D_{i}$ $:=\exp\!\left(-\left[{x_{i}-x_{i-1}}\right]\big/{\sigma}\right),$
$R_{i}$ $:=(R_{i-1}+1)\,D_{i},$
$L_{i}$ $:=(L_{i+1}+1)\,D_{i+1}.$ (49)
Then, for each $i=1,\dots,n$, the off-diagonal row sum of the Laplace kernel matrix satisfies
$\sum_{j\neq i}k_{\mathrm{L}}(x_{i},x_{j})=R_{i}+L_{i}.$ (50)
Algorithm 2 give the prefix-suffix accumulators for Laplace kernel.
Algorithm 2 Prefix-suffix accumulators for Laplace kernel
1:Ordered sample sequence $\mathbf{S}=\{s_{1},\dots,s_{l}\}$, kernel parameter $\sigma$
2:Prefix accumulators $\{R_{i}\}_{i=1}^{l}$ and suffix accumulators $\{L_{i}\}_{i=1}^{l}$
[…]
4:for $i=2$ to $l$ do
5:  $D_{i}\Leftarrow\exp\!\bigl(-(s_{i}-s_{i-1})/\sigma\bigr)$
6:  $R_{i}\Leftarrow(R_{i-1}+1)\cdot D_{i}$
7:end for
8:for i=l1i=l-1 downto 11 do
9:  $D_{i+1}\Leftarrow\exp\!\bigl(-(s_{i+1}-s_{i})/\sigma\bigr)$
10:  $L_{i}\Leftarrow(L_{i+1}+1)\cdot D_{i+1}$
11:end for
12:return $\{R_{i}\}_{i=1}^{l},\{L_{i}\}_{i=1}^{l}$
Proposition 2.
Let $x_{1}\leq\dots\leq x_{n}$ and $y_{1}\leq\dots\leq y_{m}$ be two sorted samples. Let $\mathbf{Z}=\{z_{1},\dots,z_{n+m}\}$ denote their merged sorted sequence, and let $\ell(k)\in\{\mathbf{X},\mathbf{Y}\}$ indicate the origin of $z_{k}$. Define $\Delta_{k}=z_{k}-z_{k-1}$ for $k\geq 2$, and consider the Laplacian kernel. Define the prefix accumulator $\mathcal{A}_{x\to y}(k)$ recursively by
$\mathcal{A}_{x\to y}(1)$ $=0,$ (51)
$\mathcal{A}_{x\to y}(k)$ $=\exp\!\left(-{\Delta_{k}}\big/{\sigma}\right)\mathcal{A}_{x\to y}(k-1)+\mathbb{I}\{\ell(k)=\mathbf{Y}\},$ (52)
where $\mathbb{I}\{\cdot\}$ denotes the indicator function.
Similarly, define the suffix accumulator $\mathcal{Z}_{x\to y}(k)$ by
$\mathcal{Z}_{x\to y}(n+m)$ $=0,$ (53)
$\mathcal{Z}_{x\to y}(k)$ $=\exp\!\left(-{\Delta_{k+1}}\big/{\sigma}\right)\mathcal{Z}_{x\to y}(k+1)+\mathbb{I}\{\ell(k)=\mathbf{Y}\},$ (54)
for $k=n+m-1,\dots,1$.
Then, for any $x_{i}$ corresponding to index $k$ in the merged sequence $\mathbf{Z}$, the $i$-th row sum of the cross-kernel matrix satisfies
$\sum_{j=1}^{m}k_{\mathrm{L}}(x_{i},y_{j})=\mathcal{A}_{x\to y}(k)+\mathcal{Z}_{x\to y}(k).$ (55)
The associated algorithm 3 computes the cross-kernel prefix and suffix accumulators for the sequences $\mathbf{X}$ and $\mathbf{Y}$.
Algorithm 3 Cross-kernel prefix and suffix accumulators
1:Sorted sequences $\mathbf{X}_{s}=\{x_{1},\dots,x_{n}\}$, $\mathbf{Y}_{s}=\{y_{1},\dots,y_{m}\}$; kernel parameter $\sigma$
2:Prefix and suffix accumulators: $\mathcal{A}_{x\to y},\mathcal{Z}_{x\to y},\mathcal{A}_{y\to x},\mathcal{Z}_{y\to x}$
3:Merge $\mathbf{X}_{s}$ and $\mathbf{Y}_{s}$ into sorted sequence $\mathbf{Z}=\{z_{1},\dots,z_{n+m}\}$
4:Record origin of each $z_{k}$ as $\ell(k)\in\{\mathbf{X},\mathbf{Y}\}$
5:Compute $\Delta_{k}=z_{k}-z_{k-1}$ for $k=2,\dots,n+m$
[…]
8:for $k=2$ to $n+m$ do
9:  if $\ell(k)=\mathbf{Y}$ then
10:   $\mathcal{A}_{x\to y}[k]\Leftarrow\exp(-\Delta_{k}/\sigma)\cdot\mathcal{A}_{x\to y}[k-1]+1$
11:  else
12:   $\mathcal{A}_{x\to y}[k]\Leftarrow\exp(-\Delta_{k}/\sigma)\cdot\mathcal{A}_{x\to y}[k-1]$
13:  end if
14:  if $\ell(k)=\mathbf{X}$ then
15:   $\mathcal{A}_{y\to x}[k]\Leftarrow\exp(-\Delta_{k}/\sigma)\cdot\mathcal{A}_{y\to x}[k-1]+1$
16:  else
17:   $\mathcal{A}_{y\to x}[k]\Leftarrow\exp(-\Delta_{k}/\sigma)\cdot\mathcal{A}_{y\to x}[k-1]$
18:  end if
19:end for
20:for k=n+m1k=n+m-1 downto 11 do
21:  if $\ell(k)=\mathbf{Y}$ then
22:   $\mathcal{Z}_{x\to y}[k]\Leftarrow\exp(-\Delta_{k+1}/\sigma)\cdot\mathcal{Z}_{x\to y}[k+1]+1$
23:  else
24:   $\mathcal{Z}_{x\to y}[k]\Leftarrow\exp(-\Delta_{k+1}/\sigma)\cdot\mathcal{Z}_{x\to y}[k+1]$
25:  end if
26:  if $\ell(k)=\mathbf{X}$ then
27:   $\mathcal{Z}_{y\to x}[k]\Leftarrow\exp(-\Delta_{k+1}/\sigma)\cdot\mathcal{Z}_{y\to x}[k+1]+1$
28:  else
29:   $\mathcal{Z}_{y\to x}[k]\Leftarrow\exp(-\Delta_{k+1}/\sigma)\cdot\mathcal{Z}_{y\to x}[k+1]$
30:  end if
31:end for
32:return $\mathcal{A}_{x\to y},\mathcal{Z}_{x\to y},\mathcal{A}_{y\to x},\mathcal{Z}_{y\to x}$
Remark 2.
When both sequences $\mathbf{X}$ and $\mathbf{Y}$ are given but not necessarily sorted, the algorithm first sorts each sequence individually. The computational complexity is therefore dominated by these sorting steps:
  • Sorting $\mathbf{X}$ and $\mathbf{Y}$ requires $\mathcal{O}(n\log n)$ and $\mathcal{O}(m\log m)$ operations, respectively, where $n=|\mathbf{X}|$ and $m=|\mathbf{Y}|$.
  • Once sorted, the merged sequence $\mathbf{Z}=\mathbf{X}\cup\mathbf{Y}$ can be obtained in linear time $\mathcal{O}(n+m)$.
  • The row sums of the $\mathbf{X}$-only and $\mathbf{Y}$-only kernel matrices can be computed in linear time $\mathcal{O}(n)$ and $\mathcal{O}(m)$ using recursive prefix and suffix accumulators.
  • The row sums of the cross-kernel matrix are obtained in a single forward and backward sweep over the merged sorted sequence, also requiring $\mathcal{O}(n+m)$ operations.
  • Proposition 2 and Eq. 17 handle cross terms differently by explicitly storing per-row sums, but this does not change the overall complexity.
Therefore, the overall time complexity of euMMD in this setting is $\mathcal{O}(n\log n+m\log m)$, dominated by sorting, while all subsequent accumulation and row sum computations remain linear. This is consistent with [5].
Lemma 3.
The $p$-th power of the Laplace kernel is equivalent to a Laplace kernel with bandwidth scaled by a factor of $p$. That is,
$k_{\mathrm{L}}(x,x^{\prime};\sigma)^{p}=k_{\mathrm{L}}(x,x^{\prime};\sigma/p)=\exp\left(-{p|x-x^{\prime}|}\big/{\sigma}\right)$ (56)
Proof.
By the exponential power rule, $\left[\exp\left(-{|x-x^{\prime}|}/{\sigma}\right)\right]^{p}=\exp\left(-{p|x-x^{\prime}|}/{\sigma}\right)$. ∎
4.2 Accelerated computation of variance components
We also name our method euMMD, for efficient univariate Maximum Mean Discrepancy [5], and extend it to include accelerated variance estimation. Let $x_{1}\leq\dots\leq x_{n}$ and $y_{1}\leq\dots\leq y_{m}$ be sorted samples. Define the prefix and suffix row sums for the single-sample kernel matrices as in Proposition 1:
$R_{i}^{\mathbf{X}},L_{i}^{\mathbf{X}},\,i=1,\dots,n;\,R_{j}^{\mathbf{Y}},L_{j}^{\mathbf{Y}},\,j=1,\dots,m,$
and the cross-sample prefix/suffix accumulators:
$\mathcal{A}_{x\to y}(i),\,\mathcal{Z}_{x\to y}(i),\,i=1,\dots,n;\,\mathcal{A}_{y\to x}(j),\,\mathcal{Z}_{y\to x}(j),\,j=1,\dots,m.$
4.2.1 Accelerated computation of first-order components
The unbiased estimators for each sample are
$\hat{\mathcal{U}}(x_{i})$ $=\frac{1}{n-1}\bigl(R_{i}^{\mathbf{X}}+L_{i}^{\mathbf{X}}\bigr)-\frac{1}{m}\bigl(\mathcal{A}_{x\to y}(i)+\mathcal{Z}_{x\to y}(i)\bigr),$ (57)
$\hat{\mathcal{V}}(y_{j})$ $=\frac{1}{m-1}\bigl(R_{j}^{\mathbf{Y}}+L_{j}^{\mathbf{Y}}\bigr)-\frac{1}{n}\bigl(\mathcal{A}_{y\to x}(j)+\mathcal{Z}_{y\to x}(j)\bigr).$ (58)
4.2.2 Accelerated computation of second-order components
To achieve $O(N)$ computation of the second-order terms, we leverage the bandwidth scaling property (Lemma 4.1). Specifically, while the row sums of the kernel matrix $\mathbf{K}$ are computed using the base bandwidth $\sigma$, the squared kernel terms required for the Frobenius norm $\|\mathbf{K}\|_{F}^{2}=\sum_{i,j}k(x_{i},x_{j})^{2}$ are equivalent to a Laplace kernel with bandwidth $\sigma/2$.
Therefore, we introduce a second pass of prefix-suffix accumulators with $\sigma/2$, denoted as $\{R^{\text{sq}},L^{\text{sq}}\}$ for within-sample matrices and $\{\mathcal{A}^{\text{sq}},\mathcal{Z}^{\text{sq}}\}$ for cross-sample matrices:
For the $X$ single-sample kernel matrix:
$\|\tilde{\mathbf{K}}_{XX}\|_{F}^{2}$ $=\sum_{i=1}^{n}(R_{i}^{\mathrm{sq},\mathbf{X}}+L_{i}^{\mathrm{sq},\mathbf{X}}),$ (59)
$\|\tilde{\mathbf{K}}_{XX}\mathbf{1}\|_{2}^{2}$ $=\sum_{i=1}^{n}(R_{i}^{\mathbf{X}}+L_{i}^{\mathbf{X}})^{2},$ (60)
$(\mathbf{1}^{\top}\tilde{\mathbf{K}}_{XX}\mathbf{1})^{2}$ $=\Big(\sum_{i=1}^{n}(R_{i}^{\mathbf{X}}+L_{i}^{\mathbf{X}})\Big)^{2}.$ (61)
For the $Y$ single-sample kernel matrix:
$\|\tilde{\mathbf{K}}_{YY}\|_{F}^{2}$ $=\sum_{j=1}^{m}(R_{j}^{\mathrm{sq},\mathbf{Y}}+L_{j}^{\mathrm{sq},\mathbf{Y}}),$ (62)
$\|\tilde{\mathbf{K}}_{YY}\mathbf{1}\|_{2}^{2}$ $=\sum_{j=1}^{m}(R_{j}^{\mathbf{Y}}+L_{j}^{\mathbf{Y}})^{2},$ (63)
$(\mathbf{1}^{\top}\tilde{\mathbf{K}}_{YY}\mathbf{1})^{2}$ $=\Big(\sum_{j=1}^{m}(R_{j}^{\mathbf{Y}}+L_{j}^{\mathbf{Y}})\Big)^{2}.$ (64)
For the cross-sample kernel matrix:
$\|\mathbf{K}_{XY}\|_{F}^{2}$ $=\sum_{i=1}^{n}(\mathcal{A}^{\mathrm{sq}}_{x\to y}(i)+\mathcal{Z}^{\mathrm{sq}}_{x\to y}(i)),$ (65)
$\|\mathbf{K}_{XY}\mathbf{1}_{m}\|_{2}^{2}$ $=\sum_{i=1}^{n}(\mathcal{A}_{x\to y}(i)+\mathcal{Z}_{x\to y}(i))^{2},$ (66)
$\|\mathbf{K}_{XY}^{\top}\mathbf{1}_{n}\|_{2}^{2}$ $=\sum_{j=1}^{m}(\mathcal{A}_{y\to x}(j)+\mathcal{Z}_{y\to x}(j))^{2},$ (67)
$(\mathbf{1}_{n}^{\top}\mathbf{K}_{XY}\mathbf{1}_{m})^{2}$ $=\Big(\sum_{i=1}^{n}(\mathcal{A}_{x\to y}(i)+\mathcal{Z}_{x\to y}(i))\Big)^{2}.$ (68)
Based on the theoretical analysis, Algorithm 4 presents an accelerated procedure for computing the Maximum Mean Discrepancy (MMD) with the Laplace kernel together with its full Hoeffding variance estimator, reducing the computational complexity from the naive $\mathcal{O}(N^{2})$ to $\mathcal{O}(n\log n+m\log m)$.
Algorithm 4 Fast computation of MMD2 and its variance with the Laplace kernel
1:Samples $\mathbf{X}$, $\mathbf{Y}$, parameter $\sigma$
[…]
4:Sort $\mathbf{X},\mathbf{Y}\to\mathbf{X}_{s},\mathbf{Y}_{s}$
[…]
32:return $\widehat{\mathrm{MMD}}^{2},\widehat{\operatorname{Var}}(\mathcal{T}_{1})+\widehat{\operatorname{Var}}(\mathcal{T}_{2})$
5 Numerical experiments
5.1 Variance analysis under heterogeneous settings
To evaluate the robustness and accuracy of our proposed accelerated MMD variance estimator, we conduct numerical experiments across four distinct scenarios. We consider samples drawn from Laplace distributions, i.e., $\mathbf{X}\sim\mathbb{P}=\text{Laplace}(0,1)$ and $\mathbf{Y}\sim\mathbb{Q}=\text{Laplace}(\Delta,1)$. The sample size for $\mathbf{X}$ is fixed at $n$, while we manipulate the sample size $m$ of $\mathbf{Y}$ and the distribution shift $\Delta$ to simulate both null hypothesis ($\mathcal{H}_{0}$) and alternative hypothesis ($\mathcal{H}_{1}$) conditions, as well as balanced and unbalanced sample settings.
The five experimental configurations are defined as follows: case (a): $m=n$, $\Delta=0$ (Balanced samples under $\mathcal{H}_{0}$), case (b): $m=n$, $\Delta=1$ (Balanced samples under $\mathcal{H}_{1}$), case (c): $m=1.2n$, $\Delta=0$ (Unbalanced samples under $\mathcal{H}_{0}$), case (d): $m=1.2n$, $\Delta=1$. (Unbalanced samples under $\mathcal{H}_{1}$), and Case (e): distribution shift from $\Delta=0$ to $\Delta=1$, used to examine the variance under gradually changing conditions. Figure 1 illustrates the empirical distributions of randomly drawn samples for cases (a)–(d).
Figure 1: Empirical sample distributions in different cases.
Case A: Null hypothesis with balanced samples ($m=n,\Delta=0$).
To evaluate the efficacy of the proposed accelerated MMD variance estimator under the null hypothesis, we conduct benchmarks against the classical unbiased estimators [11, 28]. The experiments utilize one-dimensional Laplace and Gaussian distributions with sample sizes $n,m\in\{126,631,3162,15849\}$.
Table 1 presents the estimated variance under the equal sample size setting. The results confirm that our proposed statistic aligns perfectly with the classical unbiased MMD. However, a significant divergence appears at the variance level. The Sutherland estimator, while theoretically unbiased, relies on a fine-grained decomposition of inner product terms, where each term is corrected independently. As shown in Table 1, this unbiasedness introduces severe numerical instability when the true variance is near zero. Specifically, for the Gaussian kernel at $n=3162$, the Sutherland estimator yields an abnormally low value ($6.170\times 10^{-10}$), a known artifact where bias-correction terms fluctuate beyond the magnitude of the raw estimate.
In contrast, our proposed method strictly follows the Hoeffding decomposition, separately estimating first- and second-order projection variances. This ensures positive semi-definiteness and superior numerical stability. Notably, our variance estimates are systematically larger than those of Gretton and Sutherland. This is because, in finite samples, empirical first-order projections retain substantial fluctuations due to cross-coupling effects among samples (Remark 1), whereas second-order components remain consistent with Gretton’s estimator.
Table 1: Comparison of MMD variance estimators under equal sample sizes ($m=n$)
Sample size ($n$) 126 631 3162 15849
Laplacian Kernel
$\widehat{\operatorname{Var}}$(MMD) (Gretton et al.) 2.145e-05 8.782e-07 3.725e-08 1.499e-09
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 3.225e-05 1.324e-06 1.154e-08 2.917e-09
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 8.456e-05 3.331e-06 8.814e-08 5.970e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 2.664e-05 1.007e-06 3.831e-08 1.527e-09
Gaussian Kernel
$\widehat{\operatorname{Var}}$(MMD) (Gretton et al.) 3.363e-05 1.337e-06 5.647e-08 2.254e-09
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 4.652e-05 1.622e-06 6.170e-10 2.862e-09
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 1.296e-04 4.761e-06 1.160e-07 7.457e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 4.232e-05 1.574e-06 5.771e-08 2.297e-09
Case B: Null hypothesis with unbalanced samples ($m\neq n,\Delta=0$).
In many real-world applications, such as anomaly detection or clinical trials, the sample sizes of the two groups are often unbalanced ($m\neq n$). Conventional MMD variance estimators, such as those proposed by [11, 28], are designed for the balanced case or rely on asymptotic assumptions that may fail when the ratio $m/n$ deviates significantly from unity.
In contrast, our proposed framework naturally accommodates unbalanced data by decoupling the first-order and second-order components according to their respective sample sizes.Table 2 evaluates the performance of our method across three unbalanced scenarios ($m=0.8n$, $m=1.5n$, and $m=2n$) using both Laplacian and Gaussian kernels. The results demonstrate that the proposed estimator provides consistent and physically meaningful variance estimates across all sample ratios and scales.
Table 2: Proposed MMD variance estimates under various unbalanced sample ratios ($m\neq n$)
Base sample size ($n$) 126 631 3162 15849
Laplacian Kernel
$\widehat{\operatorname{Var}}$(MMD) ($m=0.8n$) 6.390e-05 3.037e-06 1.172e-07 3.472e-09
$\widehat{\operatorname{Var}}$(MMD) ($m=1.5n$) 6.111e-05 2.774e-06 1.059e-07 2.060e-09
$\widehat{\operatorname{Var}}$(MMD) ($m=2n$) 3.759e-05 4.624e-07 9.301e-08 1.165e-09
Gaussian Kernel
$\widehat{\operatorname{Var}}$(MMD) ($m=0.8n$) 1.076e-04 2.991e-06 2.089e-07 4.064e-09
$\widehat{\operatorname{Var}}$(MMD) ($m=1.5n$) 9.075e-05 5.526e-06 1.833e-07 2.766e-09
$\widehat{\operatorname{Var}}$(MMD) ($m=2n$) 5.334e-05 6.046e-07 9.966e-08 1.919e-09
Case C: Alternative hypothesis with balanced samples ($m=n,\Delta\neq 0$).
To evaluate the precision of the proposed estimator under the alternative hypothesis, we introduce a location shift $\Delta=1$ between the two distributions. In this regime, the MMD statistic is non-degenerate, and the first-order term of the Hoeffding decomposition dominates the total variance. We conduct a comparative analysis against the exact unbiased estimator of [28] and the normalized asymptotic variance from [30].
Table 3 summarizes the variance estimates for Laplacian and Gaussian kernels across a wide range of sample sizes. Following the statistical definition in [30]. The results demonstrate an exceptional level of numerical consensus among the three methodologies.
Table 3: Comparison of MMD variance estimators under the alternative hypothesis ($\Delta\neq 0,m=n$)
Sample size ($n$) 126 631 3162 15849
Laplacian Kernel
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 1.049e-03 1.826e-04 3.618e-05 7.232e-06
$\widehat{\operatorname{Var}}$(MMD) (Wei et al.) 1.048e-03 1.826e-04 3.618e-05 7.232e-06
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 1.077e-03 1.838e-04 3.622e-05 7.234e-06
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 2.182e-05 8.920e-07 3.509e-08 1.411e-09
Gaussian Kernel
$\widehat{\operatorname{Var}}$(MMD) (Sutherland et al.) 1.900e-03 3.270e-04 6.555e-05 1.325e-05
$\widehat{\operatorname{Var}}$(MMD) (Wei et al.) 1.894e-03 3.268e-04 6.555e-05 1.325e-05
$\widehat{\operatorname{Var}}$(MMD) (Proposed) 1.939e-03 3.287e-04 6.563e-05 1.325e-05
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Proposed) 3.327e-05 1.386e-06 5.382e-08 2.189e-09
As shown in Table 3, for all sample sizes, our proposed variance estimates are consistently larger than those of Wei. However, when the second-order term in our estimator is subtracted, the values align with Wei’s estimates. Since Wei’s method only approximates using the first-order term, this demonstrates the higher accuracy of our approach [30].
Case D: Alternative hypothesis with unbalanced samples ($m\neq n,\Delta\neq 0$).
Then we consider the most general and challenging scenario: variance estimation under the alternative hypothesis with unbalanced sample sizes ($m\neq n$). In this regime, only the asymptotic approach of [30] and our proposed framework are applicable. We evaluate three representative sample ratios, $m=0.8n$, $m=1.5n$, and $m=2n$, to assess the consistency of the estimators.
Tables 46 report the variance estimates for the Laplacian and Gaussian kernels. Across all sample ratios, our proposed estimator exhibits excellent agreement with the normalized results of [30]. As the sample size $n$ increases, both methods consistently recover the expected decay rate of $\mathcal{O}(m^{-1}+n^{-1})$.
In addition, our framework provides an explicit estimate of the second-order variance component. Consistent with the earlier discussion, since the approach of [30] effectively removes the first-order projection, our full variance estimates are systematically larger. Once the second-order term is subtracted, the remaining variance closely matches Wei et al.’s estimator, further confirming the correctness and finer granularity of our decomposition.
Table 4: Comparison of MMD variance estimators under unbalanced alternative hypothesis ($\Delta=1,m=0.8n$).
Kernel Estimator $n=126$ $n=631$ $n=3162$ $n=15849$
Laplacian Wei et al. 1.592e-03 2.315e-04 4.535e-05 8.358e-06
Proposed 1.627e-03 2.330e-04 4.541e-05 8.361e-06
Proposed ($\widehat{\operatorname{Var}}(\mathcal{T}_{2})$) 2.306e-05 1.029e-06 4.100e-08 1.657e-09
Gaussian Wei et al. 2.947e-03 4.027e-04 8.350e-05 1.516e-05
Proposed 3.000e-03 4.050e-04 8.359e-05 1.516e-05
Proposed ($\widehat{\operatorname{Var}}(\mathcal{T}_{2})$) 3.274e-05 1.576e-06 6.254e-08 2.529e-09
Table 5: Comparison of MMD variance estimators under unbalanced alternative hypothesis ($\Delta=1,m=1.5n$).
Kernel Estimator $n=126$ $n=631$ $n=3162$ $n=15849$
Laplacian Wei et al. 1.021e-03 1.656e-04 2.728e-05 6.167e-06
Proposed 1.037e-03 1.663e-04 2.730e-05 6.168e-06
Proposed ($\widehat{\operatorname{Var}}(\mathcal{T}_{2})$) 1.066e-05 4.931e-07 1.950e-08 7.640e-10
Gaussian Wei et al. 1.858e-03 3.083e-04 4.915e-05 1.120e-05
Proposed 1.883e-03 3.094e-04 4.919e-05 1.120e-05
Proposed ($\widehat{\operatorname{Var}}(\mathcal{T}_{2})$) 1.500e-05 7.369e-07 2.888e-08 1.119e-09
Table 6: Comparison of MMD variance estimators under unbalanced alternative hypothesis ($\Delta=1,m=2n$).
Kernel Estimator $n=126$ $n=631$ $n=3162$ $n=15849$
Laplacian Wei et al. 9.851e-04 1.294e-04 2.796e-05 5.550e-06
Proposed 9.966e-04 1.298e-04 2.797e-05 5.550e-06
Proposed ($\widehat{\operatorname{Var}}(\mathcal{T}_{2})$) 5.169e-06 2.130e-07 8.078e-09 3.180e-10
Gaussian Wei et al. 1.786e-03 2.399e-04 5.093e-05 1.016e-05
Proposed 1.803e-03 2.404e-04 5.095e-05 1.016e-05
Proposed ($\widehat{\operatorname{Var}}(\mathcal{T}_{2})$) 6.238e-06 2.633e-07 9.289e-09 3.605e-10
Case E: Effect of increasing distribution shift on the variance.
As the distribution shift parameter $\delta$ increases, the test statistic undergoes a gradual transition from the null hypothesis to the alternative hypothesis. In our experiments, for each configuration we fix the sample sizes and the underlying sample realizations, and vary only the shift parameter $\delta$ by applying a deterministic translation to one sample set. This experimental design isolates the effect of distributional shift and ensures that the observed variance evolution is not confounded by additional sampling variability. Figure 2 illustrates the estimated total variance and the second-order variance component under different sample size configurations $n=800,,m\in{800,1500,2000}$}.
Figure 2: Growth trend of the variance under increasing distribution shift.
Figure 2(a) shows that the total variance increases smoothly and continuously as $\delta$ grows, while Figure 2(b) depicts the evolution of the second-order variance component. This behavior is consistent with the theoretical analysis of [11]: under the null hypothesis, the MMD statistic is dominated by the second-order (degenerate) component, whereas under the alternative hypothesis the first-order projection becomes dominant. However, our proposed unified variance estimator, by employing empirical projections for the first-order terms, ensures a continuous transition across the null–alternative boundary, a property that is particularly desirable in modern machine learning settings where stability under small distributional perturbations is crucial.
5.2 Accuracy and efficiency of Laplacian-based accelerated euMMD
Statistical accuracy of euMMD.
To validate the statistical precision of the proposed accelerated euMMD estimator, we conduct a comparative analysis under the alternative hypothesis $\mathcal{H}_{1}$. We consider one-dimensional Laplace distributions where $\mathbb{P}\sim\mathrm{Laplace}(0,1)$ and $\mathbb{Q}\sim\mathrm{Laplace}(1,1)$, representing a location shift. To assess the performance under imbalanced conditions, the sample sizes are set at a ratio of $m=1.2n$, with $n\in\{10,100,1000,10000\}$.
As shown in Table 7, the proposed euMMD statistic and its variance components ($\mathcal{T}_{1}$ and $\mathcal{T}_{2}$) are algebraically identical to the classical matrix-form unbiased MMD estimator. This alignment confirms that our $O(N\log N)$ framework is not an approximation but an exact acceleration of the finite-sample U-statistic.
Notably, in the small sample regime ($n=10$), the estimator maintains rigorous consistency. By carefully aligning the prefix-suffix accumulators to exclude diagonal self-similarity terms $k(x_{i},x_{i})$, the resulting second-order component $\mathcal{T}_{2}$ remains positive and numerically precise. As $n$ increases, the variance estimates exhibit the expected $O(n^{-1})$ decay, with the first-order component $\mathcal{T}_{1}$ dominating the total variance. These results demonstrate that euMMD preserves the full statistical fidelity of the MMD while significantly enhancing computational efficiency for large-scale data analysis.
Table 7: Numerical Consistency Verification: Matrix-form MMD vs. Accelerated euMMD
Sample size ($n,m$) (10, 12) (100, 120) (1000, 1200) (10000, 12000)
$\widehat{\mathrm{MMD}}^{2}$ (Matrix) 5.206e-02 1.272e-01 1.193e-01 1.030e-01
$\widehat{\mathrm{MMD}}^{2}$ (euMMD) 5.206e-02 1.272e-01 1.193e-01 1.030e-01
$\widehat{\operatorname{Var}}_{\text{total}}$ (Matrix) 1.465e-02 1.368e-03 1.202e-04 1.045e-05
$\widehat{\operatorname{Var}}_{\text{total}}$ (euMMD) 1.465e-02 1.368e-03 1.202e-04 1.045e-05
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (Matrix) 3.355e-03 2.772e-05 2.740e-07 2.843e-09
$\widehat{\operatorname{Var}}$($\mathcal{T}_{2}$) (euMMD) 3.355e-03 2.772e-05 2.740e-07 2.843e-09
Note: All results are identical up to machine precision, validating the exactness of the $O(N\log N)$ implementation.
Efficiency of variance estimation
To evaluate the computational efficiency and scalability of the proposed accelerated framework for large-scale variance estimation, we benchmark its runtime and memory performance against the traditional $O(n^{2})$ algorithm. Due to the quadratic explosion in space complexity, the traditional method is restricted to sample sizes $n\leq 10,000$, beyond which it encounters hardware memory limitations. In contrast, leveraging the near-linear complexity and extreme memory efficiency of our algorithm, we extend the sample size to $n=10^{6}$. All experiments are independently repeated 20 times, with average results reported in Figure 3.
Figure 3: Double-logarithmic comparison of runtime and memory consumption for both algorithms under unequal sample sizes
Under the double-logarithmic scale, the performance divergence between the two paradigms is striking. In terms of computational time, the traditional algorithm exhibits a rigid quadratic growth pattern ($\mathcal{O}(n^{2})$). Conversely, the proposed accelerated algorithm demonstrates a remarkably mild growth rate that aligns with its theoretical $\mathcal{O}(n\log n)$ complexity. Remarkably, even at the million-sample scale ($n=10^{6}$), our algorithm completes the variance estimation within seconds, whereas the traditional approach would theoretically require hours of computation, assuming infinite memory.
The contrast in memory consumption is even more pronounced. The kernel matrix method suffers from a rapid expansion in space usage; at $n=10,000$, its peak memory footprint exceeds 3 GB, rendering larger-scale scenarios infeasible due to Out-Of-Memory (OOM) errors. In contrast, the accelerated algorithm maintains an almost flat trajectory in the log-log plots. Its memory footprint remains consistently below the 1 MB threshold across all tested scales—essentially negligible compared to system background usage. This "near-zero" space overhead confirms that our method successfully shifts the bottleneck of MMD variance estimation from hardware capacity to basic CPU cycles.
6 Conclusion
In this work, we studied the variance of the Maximum Mean Discrepancy (MMD) statistic from the perspective of its U-statistic representation and Hoeffding decomposition. Our contributions are threefold:
  • Unified finite-sample variance characterization: We established analytical expressions for the MMD variance that hold across different hypotheses (null and alternative) and sample configurations (balanced and unbalanced). This provides a consistent framework for understanding variance behavior in practical two-sample testing scenarios.
  • Exact acceleration for univariate MMD: For the Laplace kernel in the univariate case, we developed an exact, quasi-linear-time algorithm that reduces the computational complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$ while preserving the full Hoeffding variance estimator. This allows efficient, accurate variance computation even for large sample sizes.
Overall, the proposed framework builds a bridge between theoretical understanding and practical computation, making variance estimation and inference in finite-sample two-sample testing more robust and efficient.
Appendix A Auxiliary Lemmas for accelerated computation with the Laplacian kernel
Lemma 1.
Let $\{z_{1},z_{2},\dots,z_{m}\}\subset\mathbb{R}$ and let $f:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a symmetric function, i.e., $f(x,y)=f(y,x)$. Then
$\sum_{i=1}^{m}\sum_{j=1}^{i}f(z_{i},z_{j})=\sum_{i=1}^{m}\sum_{j=i}^{m}f(z_{i},z_{j}).$
Lemma 2.
Under the conditions of Lemma 1, we have
$\sum_{i=1}^{m}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}f(z_{i},z_{j})=2\sum_{i=1}^{m}\sum_{j=1}^{i-1}f(z_{i},z_{j}).$
Lemma 3.
Let $\beta:\{1,2,\dots,m\}\to\{1,2,\dots,m\}$ be any permutation. Then
$\sum_{i=1}^{m}\sum_{j=1}^{i-1}f(z_{i},z_{j})=\sum_{i=1}^{m}\sum_{j=1}^{i-1}f\!\left(z_{\beta(i)},z_{\beta(j)}\right).$
Proposition 4.
Let $\{z_{1},\dots,z_{m}\}\subset\mathbb{R}$ and denote the order statistics by $z_{(1)}\leq z_{(2)}\leq\cdots\leq z_{(m)}$. Consider the Laplacian kernel with bandwidth $\sigma>0$, $k_{\mathrm{L}}(x,y)=\exp\!\left(-{|x-y|}\big/{\sigma}\right)$. Define
$S:=\sum_{i=1}^{m}\sum_{j=1}^{i-1}k_{\mathrm{L}}(z_{i},z_{j})=\sum_{i=1}^{m}\sum_{j=1}^{i-1}\exp\!\left(-[{z_{(i)}-z_{(j)}}]\big/{\sigma}\right).$
Then $S$ can be computed recursively as follows. Set $S_{1}=R_{1}=0$, and for $k=2,\dots,m$, define
$D_{k}$ $=\exp\!\left(-[{z_{(k)}-z_{(k-1)}}]\big/{\sigma}\right),$
$R_{k}$ $=(R_{k-1}+1)\,D_{k},$
$S_{k}$ $=S_{k-1}+R_{k}.$
Then $S=S_{m}$.